### v2

# DeformTime: Capturing Variable Dependencies with Deformable Attention for Time Series Forecasting

Abstract
In multivariate time series (MTS) forecasting, existing state-of-the-art deep learning approaches tend to focus on autoregressive formulations and overlook the information within exogenous indicators. To address this limitation, we present DeformTime, [^1] The source code of DeformTime is available at github.com/ClaudiaShu/DeformTime. a neural network architecture that attempts to capture correlated temporal patterns from the input space, and hence, improve forecasting accuracy. It deploys two core operations performed by deformable attention blocks (DABs): learning dependencies across variables from different time steps (variable DAB), and preserving temporal dependencies in data from previous time steps (temporal DAB). Input data transformation is explicitly designed to enhance learning from the deformed series of information while passing through a DAB. We conduct extensive experiments on $6$ MTS data sets, using previously established benchmarks as well as challenging infectious disease modelling tasks with more exogenous variables. The results demonstrate that DeformTime improves accuracy against previous competitive methods across the vast majority of MTS forecasting tasks, reducing the mean absolute error by $10\%$ on average. Notably, performance gains remain consistent across longer forecasting horizons.
1 Introduction
Time series forecasting models provide multifaceted solutions for many application domains, including health [39, 15], climate [37], energy [7], transport [24], and finance [38]. In their pursuit for greater accuracy, forecasting methods have always been trying to benefit from multimodal exogenous predictors [6, 1]. Over recent years, a plethora of new digitised information sources has become available, accompanied by a rapid development of capable and efficient deep learning architectures [35, 19]. However, current state-of-the-art (SOTA) models for multivariate time series (MTS) forecasting favour the deployment of many-to-many sequence deep learning architectures [32, 46, 29, 44]. We argue that this modelling approach does not necessarily provide the right learning mechanism for tasks with a definitive target (endogenous) variable and a considerable amount of exogenous variables. Furthermore, although various neural networks (NNs) were proposed to address MTS forecasting challenges [55, 51, 13, 29], many [53, 32, 16, 23, 17] were not designed to incorporate inter-variate dependencies, a key property of conventional forecasters [34, 14, 3].
There certainly exist neural MTS forecasting models that attempt to leverage dependencies across variables and time. Solutions based on recurrent neural network (RNN) architectures [19, 25] can capture information from previous time steps (up to an extent), but have a limited capacity in establishing inter-variate dependencies. LightTS [54] uses multi-layer perceptrons to model dependencies within the input data, first along the temporal dimension and then across variables. For transformer-based models, Crossformer [55] proposes to first partition the input into temporal patches, and then use a router module to capture information across variables from different time steps, improving performance compared to pre-established RNN and Graph NN architectures [19, 47]. However, both LightTS and Crossformer have been outperformed by later proposed models that do not establish dependencies between variables [53, 32, 16, 23, 17]. The counterargument from these approaches is that forecasters without inter-variable dependency modules benefit from being able to use longer look-back windows without significantly increasing model complexity [11]. Nevertheless, a common issue throughout the deep learning MTS forecasting literature that casts doubt on some of these conclusions is the inappropriateness of various benchmark tasks (see Appendix A.5). In addition, recent work has argued that to improve performance while exploiting inter-variable dependencies, more effort is required in temporally aligning the input time series [56]. Motivated by this, we have introduced guided re-arrangements of the input to better capture inter- and intra-variable dynamics.
Deformable neural networks were initially proposed in computer vision [5] to accommodate geometric transformations with Convolutional Neural Networks (CNNs). Combined with transformer-based NN architectures, deformation has achieved SOTA performance in various tasks [60, 4, 48]. Combining our aforementioned insights with the concept of deformation has led to the development of DeformTime, a novel MTS forecasting model that deploys a deformable module on top of transformer encoders to introduce some flexibility in the determination of receptive fields across different variables and time steps. The premise of DeformTime is the inclusion of learnable mechanisms that we refer to as deformable attention blocks (DABs). These enable transformations of the input information stream that enhance learning from key patterns within endo- but most importantly exogenous variables, an operation that ultimately improves forecasting accuracy.
We summarise the main contributions of this paper as follows:
  1. (a)
    We propose DeformTime, a novel MTS forecasting model, that better captures inter- and intra-variate dependencies at different temporal granularities. It comprises two DABs which facilitate learning from adaptively transformed input across variables (V-DAB) and time (T-DAB).
  2. (b)
    We assess the predictive accuracy of MTS forecasters on $3$ established benchmarks as well as $3$ new infectious disease prevalence tasks, each across $4$ increasing horizons. We argue that the disease prevalence tasks support a more thorough evaluation because data sets include a substantial amount of exogenous variables, and encompass multiple years, $4$ of which are used as distinct annual test periods.
  3. (c)
    In our experiments, DeformTime reduces the mean absolute error (MAE) by $10\%$ on average compared to the most competitive forecasting method (can be a different one each time). There is a $7.4\%$ MAE reduction based on established benchmarks and $12.6\%$ for the disease forecasting tasks. Overall, performance gains remain stable as the forecasting horizon increases.
  4. (d)
    As an aside, we note that models that attempt to establish some form of inter-variable dependency performed better in evidently (based on the results) the more challenging task of disease forecasting. This also highlights the need for more appropriate methods of assessment.
2 MTS forecasting task definition
Across our experiments, we consider an MTS forecasting task where the focus is on a single target variable, i.e. there might exist multiple inputs, and optionally multiple outputs, but we are only concerned with the predictions of one output variable. All models follow a rolling window forecasting setting with a look-back window of $L$ time steps. To be more precise, at time step $t$, $\mathbf{Q}_{t}\!\in\!\mathbb{R}^{L\times C}$ holds the time series of $C$ exogenous variables over the $L$ time steps $\{t\!-\!L\!+\!1,\dots,t\!-\!1,t\}$. In addition, $\mathbf{y}_{t-\delta}\!\in\!\mathbb{R}^{L}$ holds the corresponding $L$ autoregressive historical values for the target variable for time steps $\{t\!-\!\delta\!-\!L\!+\!1,\dots,t\!-\!\delta\!-\!1,t\!-\!\delta\}$. Note that $\delta\!\in\!\mathbb{N}_{0}$ introduces an optional delay of $\delta$ time steps between the observed exogenous variables and the endogenous (target) variable. This becomes relevant in a real-time infectious disease forecasting task, where estimates for the rate of an infectious disease are becoming available with a delay of $\delta=7$ or $14$ days, when other indicators are not affected by this (see Appendix A.2). Joining both input signals, we obtain $\mathbf{Z}_{t}=\left[\mathbf{Q}_{t},\mathbf{y}_{t-\delta}\right]\!\in\!\mathbb{R}^{L\times\left(C+1\right)}$. Commonly, the prediction target is denoted by $\mathbf{y}_{t+H}\!\in\!\mathbb{R}^{H}$, where $H$ indicates the number of time steps we are forecasting ahead. However, some (baseline) models follow a multi-task learning formulation whereby every input variable becomes a prediction target. On this occasion, the prediction targets are denoted by $\mathbf{Y}_{t+H}=\left[\mathbf{Q}_{t+H},\mathbf{y}_{t+H}\right]\!\in\!\mathbb{R}^{H\times\left(C+1\right)}$. Hence, the aim of the forecasting task is to learn a function $f\!\left(\cdot\right)$ such that $f\!:\mathbf{Z}_{t}\rightarrow\mathbf{y}_{t+H}$ or $\mathbf{Y}_{t+H}$. Irrespective of the number of outputs, forecasting accuracy is reported on the final estimate of $\mathbf{y}_{t+H}$, $y_{t+H}\!\in\!\mathbb{R}$, which denotes the value of the target variable at time step $t\!+\!H$. For notational simplicity, we have chosen to only imply temporal subscripts for the remainder of the manuscript (i.e. use $\mathbf{Z}$ over $\mathbf{Z}_{t}$).
Figure 1: The architecture of DeformTime. We use the notation introduced in sections 2 and 3. DeformTime’s core modules comprise two deformable attention blocks (DABs), a variable DAB (V-DAB) and a temporal DAB (T-DAB) that respectively capture inter- and intra-variable dependencies. Both DABs reside in the Encoder module. We deploy a 2-layer GRU as the Decoder. Finally, we have visualised key data operations (Segment and Adapt blocks) that take place in the DABs.
3 Time series forecasting with DeformTime
An overview of DeformTime’s structure is presented in Figure 1. Motivated by our assumption that variables (endo- and exogenous) in an MTS task may not only be correlated with the historical values of the target variable, but also with other predictors within adjacent time steps, DeformTime deploys a modified transformer encoder layer with two core modules that attempt to account for both dependencies: a Variable Deformable Attention Block (V-DAB) to capture inter-variable dependencies across time, and a Temporal Deformable Attention Block (T-DAB) to capture intra-variable dependencies across different time periods. In the following sections, we provide a detailed description of all modules and operations of DeformTime.
3.1 Multi-head attention
We first revisit the multi-head attention mechanism originally proposed by Vaswani et al. [41]. Given an input $\mathbf{Z}\!\in\!\mathbb{R}^{v\times u}$, $\mathbf{Q},\mathbf{K},$ and $\mathbf{V}\!\in\!\mathbb{R}^{v\times u}$ respectively denote the query, key, and value embeddings projected from $\mathbf{Z}$ with learnable linear projection matrices $\mathbf{W}_{Q},\mathbf{W}_{K},$ and $\mathbf{W}_{V}\!\in\!\mathbb{R}^{u\times u}$ (as in $\mathbf{Q}=\mathbf{Z}\mathbf{W}_{Q}$). Assuming $M$ attention heads, these embeddings are each partitioned into $M$ non-overlapping submatrices column-wise. For example, the $m$-th submatrix of $\mathbf{Q}$, denoted by $\mathbf{Q}_{m}$, has a dimensionality of $v\times(u/M)$. Following up on this notational convention, the output of the $m$-th attention head, $\mathbf{A}_{m}\!\in\!\mathbb{R}^{v\times(u/M)}$, is given by
$\mathbf{A}_{m}=\texttt{softmax}\!\left(\mathbf{Q}_{m}\mathbf{K}_{m}^{\top}/\sqrt{u/M}\right)\mathbf{V}_{m}\,.$ (1)
The outputs from each head are then concatenated along the second dimension and linearly projected into the hidden dimension $u$ with a weight matrix $\mathbf{W}\!\in\!\mathbb{R}^{u\times u}$ to form the output $\mathbf{A}\!\in\!\mathbb{R}^{v\times u}$. When we set $M=1$, we refer to this method as single-head attention.
3.2 Neighbourhood-aware input embedding
The receptive field of DeformTime is determined by a CNN with sampling offset (see section 3.3). This introduces a limitation in capturing relationships between neighbouring variables. We seek to minimise the impact of this by learning a Neighbourhood-Aware input Embedding (NAE) as follows. We first re-arrange the order of variables in the input $\mathbf{Z}$, ranking them based on their linear correlation with the target variable in a temporally aligned fashion (see also Appendix C.1). We then embed the correlation-driven re-arranged version of the input that captures $C+1$ variables to a hidden dimension $d$. However, we do not perform this embedding as a single holistic step, but embed $G$ neighbouring groups of variables, each one to a shortened embedding of size $d/G$, using a fully connected layer. [^2] The input is zero-padded to a proper length, if the number of variables is not divisible by $G$. Both $d$ and $G$ are hyperparameters that we learn during training. The embeddings of each input variable grouping are concatenated to form $\mathbf{E}\!\in\!\mathbb{R}^{L\times d}$. Finally, we also add a sinusoidal position embedding [41, 57], $\mathbf{P}_{\texttt{n}}\!\in\!\mathbb{R}^{L\times d}$, to $\mathbf{E}$, and normalise within layer as widely adopted in prior work [8, 55]. Hence, the embedding $\mathbf{Z}_{\texttt{e}}\!\in\!\mathbb{R}^{L\times d}$ of the input $\mathbf{Z}$ is given by
$\mathbf{Z}_{\texttt{e}}=\texttt{LN}\!\left(\mathbf{E}+\mathbf{P}_{\texttt{n}}\right)\,,$ (2)
where $\texttt{LN}(\cdot)$ denotes layer normalisation [2].
3.3 Variable deformable attention block (V-DAB)
[…]
To determine the offset $\Delta\mathbf{p}\!\in\!\mathbb{R}^{2}$ we conduct the following series of operations. We first obtain a query embedding, $\mathbf{Q}_{\texttt{p}}\!\in\!\mathbb{R}^{\ell\times d}$, over a patch of the input as described in section 3.1, i.e. $\mathbf{Q}_{\texttt{p}}=\mathbf{Z}_{\texttt{p}}\mathbf{W}_{Q}$. We then pass $\mathbf{Q}_{\texttt{p}}$ from a 2D CNN, denoted by $\theta_{\texttt{off}}$, comprising a convolutional layer with a $k\!\times\!k$ kernel that captures neighbouring information, and a $1\!\times\!1$ convolutional layer that projects the embedded feature. We use the $\tanh$ activation function and a learnable amplitude $\alpha\!\in\!\mathbb{R}_{>0}$ (that controls the upper bound of the sampling range) to obtain $\Delta\mathbf{p}$. Hence, $\Delta\mathbf{p}$ is given by
$\Delta\mathbf{p}=\alpha\cdot\tanh\!\left(\theta_{\texttt{off}}\!\left(\mathbf{Q}_{\texttt{p}}\right)\right)\,.$ (3)
We use $\Delta\mathbf{p}$ and bilinear interpolation ($\psi$) over a $2\!\times\!2$ grid determined by $\mathbf{p}+\Delta\mathbf{p}$ [5], to obtain the deformed patch $\mathbf{Z}_{\texttt{d}}$:
$\mathbf{Z}_{\texttt{d}}=\psi\left(\mathbf{Z}_{\texttt{p}};\mathbf{p}+\Delta\mathbf{p}\right)\,.$ (4)
We then obtain $\mathbf{K}_{\texttt{d}},\mathbf{V}_{\texttt{d}}$ from $\mathbf{Z}_{\texttt{d}}$ with
$\mathbf{K}_{\texttt{d}}=\mathbf{Z}_{\texttt{d}}\mathbf{W}_{K}\,\,\,\,\text{and}\,\,\,\,\mathbf{V}_{\texttt{d}}=\mathbf{Z}_{\texttt{d}}\mathbf{W}_{V}+\mathbf{P}_{\texttt{v}}\,,$ (5)
where $\mathbf{P}_{\texttt{v}}\!\in\!\mathbb{R}^{\ell\times d}$ is a relative positional bias [40] that is added to assist in maintaining some pairwise positional information after the deformation.
A single head self-attention layer is applied to attend $\mathbf{Q}_{\texttt{p}}$ to $\mathbf{K}_{\texttt{d}},\mathbf{V}_{\texttt{d}}$ given by
$\mathbf{A}_{i}=\left(\texttt{softmax}\!\left(\mathbf{Q}_{\texttt{p}}\mathbf{K}_{\texttt{d}}^{\top}/\sqrt{d}\right)\mathbf{V}_{\texttt{d}}\right)\mathbf{W}_{i}\,,$ (6)
where $\mathbf{W}_{i}\!\in\!\mathbb{R}^{d\times d}$ is the weight matrix, to form the attention embedding $\mathbf{A}_{i}\!\in\!\mathbb{R}^{\ell\times d}$ for the $i$-th segmented patch. The output from all patches is concatenated across the temporal dimension, forming $\mathbf{A}\!\in\!\mathbb{R}^{(n\cdot\ell)\times d}$. The final output of the V-DAB module, $\mathbf{Z}_{\texttt{v}}\!\in\!\mathbb{R}^{L\times d}$, is given by
$\mathbf{Z}_{\texttt{v}}=\mathbf{W}_{\texttt{v}}^{\top}\mathbf{A}\,,$ (7)
where $\mathbf{W}_{\texttt{v}}\!\in\!\mathbb{R}^{(n\cdot\ell)\times L}$ denotes the weight matrix of a fully connected layer.
3.4 Temporal deformable attention block (T-DAB)
[…]
The factor $r$ (time steps) can be set to different values across encoder layers to capture dependencies at various temporal granularities (see Appendix C). If we set $r=1$, $\mathbf{Z}_{\texttt{e}}$ remains unchanged. Similarly to V-DAB, we obtain the query embedding $\mathbf{Q}_{\texttt{r}}\!\in\!\mathbb{R}^{\kappa\times d}$ over the transformed input using $\mathbf{Q}_{\texttt{r}}=\mathbf{Z}_{\texttt{r}}\mathbf{U}_{Q}$, where $\mathbf{U}_{Q}\!\in\!\mathbb{R}^{d\times d}$ is a weight matrix. We then implement an augmented Transformer attention that contains deformed information across the temporal dimension. We work our way column-wise, and for every column $\mathbf{z}_{\texttt{r}}\!\in\!\mathbf{Z}_{\texttt{r}}$, we sample (linearly as opposed to the bilinear approach in V-DAB) from an index position $p+\Delta p\!\in\!\mathbb{R}$.
The position offset ($\Delta p$) is shared among groups of correlated embedded sequences as they may have similar temporal dependencies. Given that the original input variables were already grouped based on their correlation with the target signal (section 3.2), we expect that correlations are already captured in feature neighbourhoods of $\mathbf{Z}_{\texttt{r}}$. Therefore, to obtain $\Delta p$, we first divide $\mathbf{Q}_{\texttt{r}}$ into the same $G$ subgroups (column-wise) as in the NAE operation. Each subgroup, $\mathbf{Q}_{g}\!\in\!\mathbb{R}^{\kappa\times(d/G)}$ with $g=\{1,\dots,G\}$, shares the same temporal deformation controlled by $\Delta p^{(g)}$. Similarly to Eq. 3, $\Delta p^{(g)}$ is defined as
$\Delta p^{(g)}=\alpha\cdot\tanh\!\left(\eta_{\texttt{off}}\!\left(\mathbf{Q}_{g}\right)\right)\,,$ (8)
with $\eta_{\texttt{off}}$ denoting a 1D CNN (with a $k\!\times\!1$ and $1\!\times\!1$ convolution layers), and $\alpha$ being the same offset amplitude as in V-DAB. We then obtain the matrix of temporally deformed sequences $\mathbf{Z}_{\texttt{s}}=\left[\mathbf{Z}_{\texttt{s}}^{(1)},\dots,\mathbf{Z}_{\texttt{s}}^{(G)}\right]\!\in\!\mathbb{R}^{\kappa\times d}$; each submatrix, $\mathbf{Z}_{\texttt{s}}^{(g)}\!\in\!\mathbb{R}^{\kappa\times(d/G)}$, is given by
$\mathbf{Z}_{\texttt{s}}^{(g)}=\phi\!\left(\mathbf{z}_{\texttt{r}}\!\in\!\mathbf{Z}_{\texttt{r}}^{(g)};p+\Delta p^{(g)}\right)\,,$ (9)
where $\phi(\cdot)$ denotes a linear interpolation function over $2$ adjacent points of $\mathbf{z}_{\texttt{r}}\!\in\!\mathbf{Z}_{\texttt{r}}^{(g)}$ determined by the deformed index $p+\Delta p^{(g)}$.
Similarly to V-DAB, $\mathbf{K}_{\texttt{s}},\mathbf{V}_{\texttt{s}}\!\in\!\mathbb{R}^{\kappa\times d}$ are obtained from $\mathbf{Z}_{\texttt{s}}$, using $\mathbf{K}_{\texttt{s}}=\mathbf{Z}_{\texttt{s}}\mathbf{U}_{K}$, and $\mathbf{V}_{\texttt{s}}=\mathbf{Z}_{\texttt{s}}\mathbf{U}_{V}+\mathbf{P}_{\texttt{t}}$, where $\mathbf{P}_{\texttt{t}}\!\in\!\mathbb{R}^{\kappa\times d}$ is a relative positional bias. We then apply multi-head attention, by first splitting $\mathbf{K}_{\texttt{s}}$ and $\mathbf{V}_{\texttt{s}}$ into $G$ groups (submatrices) denoted by $\mathbf{K}_{g}$ and $\mathbf{V}_{g}\!\in\!\mathbb{R}^{\kappa\times(d/G)}$, and then by using $\mathbf{A}_{g}=\texttt{softmax}\!\left(\mathbf{Q}_{g}\mathbf{K}_{g}^{\top}/\sqrt{d/G}\right)\mathbf{V}_{g}$. The outputs from different heads are concatenated column-wise and linearly projected with a weight matrix $\mathbf{W}_{i}\!\in\!\mathbb{R}^{d\times d}$ to obtain the attention embedding of the $i$-th patch $\mathbf{A}_{i}\!\in\!\mathbb{R}^{\kappa\times d}$. To form the output of T-DAB, denoted by $\mathbf{Z}_{\texttt{t}}\!\in\!\mathbb{R}^{L\times d}$, the outputs of all patches are concatenated along the first dimension and re-arranged back to the original temporal structure (matching the input $\mathbf{Z}_{\texttt{e}}$ as well as V-DAB’s output $\mathbf{Z}_{\texttt{v}}$).
3.5 Encoder
The encoder of DeformTime may consist of more than one layer. In our experiments, we deploy $2$ encoder layers, similarly to related work [28, 46]. Each encoder layer contains two DAB branches (on the left and right side, respectively – see also Figure 1), comprising stacked transformer blocks [8] with layer normalisation placed within the residual connection [49, 48]. Instead of using vanilla attention, we are introducing a V-DAB (left side) and a T-DAB (right side) module to capture inter- and intra-variate dependencies, respectively. The operations of a branch can be summarised by
$\mathbf{Z}_{\texttt{i}}=\texttt{Drop}\!\left(\texttt{DAB}\left(\mathbf{Z}_{\texttt{e}}\right)\right)+\mathbf{Z}_{\texttt{e}}\,\,\text{followed by}\,\,\mathbf{Z}_{\texttt{c}}=\texttt{Drop}\!\left(\texttt{MLP}\!\left(\texttt{LN}\!\left(\mathbf{Z}_{\texttt{i}}\right)\right)\right)+\mathbf{Z}_{\texttt{e}}\,,$ (10)
where $\texttt{DAB}(\cdot)$ denotes the V-DAB or T-DAB operation, $\mathbf{Z}_{\texttt{i}},\mathbf{Z}_{\texttt{c}}\!\in\!\mathbb{R}^{L\times d}$ are an intermediate output and the output of the DAB encoder layer, respectively, $\texttt{Drop}(\cdot)$ denotes stochastic depth [12] where layers within the network are randomly dropped during training with a learnable probability, and $\texttt{MLP}(\cdot)$ is a $2$-layer perceptron with ReLU activation and a hidden size of $d$. The outputs from the two encoder branches are concatenated over the second dimension. We feed the concatenated output into a fully connected layer to form the output $\mathbf{Z}_{j}\!\in\!\mathbb{R}^{L\times d}$ of the $j$-th encoder layer.
3.6 Encoder-decoder structure
In T-DAB (section 3.4), we set the time window $r$ to different values across encoder layers, enabling the model to attend to information at multiple temporal granularities. To effectively learn from this encoding, we construct a hierarchical encoder structure motivated by prior related work [57, 26, 55]. Specifically, the output from encoder layer $j\!-\!1$, $\mathbf{Z}_{j-1}\!\in\!\mathbb{R}^{L\times d}$, becomes the input to encoder layer $j$. This is formulated by
$\mathbf{Z}_{j}=\begin{cases}\texttt{Enc}\!\left(\mathbf{Z}_{\texttt{c}}\right)&\text{if }j\!=\!1\\ \texttt{Enc}\!\left(\mathbf{Z}_{j-1}\right)&\text{otherwise}\end{cases}\,,$ (11)
where $\texttt{Enc}(\cdot)$ denotes the operations of an encoder layer with deformable attention (section 3.5). The output of the final encoder layer is fed to a $2$-layer Gated Recurrent Unit (GRU) NN with a hidden dimension of $d$ that acts as the decoder. The output of the GRU decoder, $\mathbf{Z}_{\texttt{out}}$, maintains the same dimensionality as the input, i.e. ${L\!\times\!d}$. Finally, we use a $2$-layer perceptron with a hidden dimension of $d$ and LeakyReLU activation to project $\mathbf{Z}_{\texttt{out}}$ along the temporal dimension, forming an intermediate output $\mathbf{Z}^{\prime}_{\texttt{out}}\!\in\!\mathbb{R}^{H\times d}$, where $H$ denotes the number of steps we are forecasting ahead as defined in section 2. We then feed $\mathbf{Z}^{\prime}_{\texttt{out}}$ to a fully connected layer to linearly project it into the target forecasting output $\hat{\mathbf{y}}\!\in\!\mathbb{R}^{H}$.
4 Results
We present forecasting accuracy results across $3$ benchmark and $3$ disease rate prediction tasks, comparing DeformTime to other competitive ‘baseline’ models. In addition, we provide an ablation study and a computational complexity / efficiency analysis.
4.1 Forecasting tasks, baseline methods, and experiment settings
Experiments are conducted on $6$ real-world data sets. These include $3$ established benchmarks from previously published papers [45, 53, 52, 46, 32, 29], namely a weather and $2$ electricity transformer temperature (ETTh1 and ETTh2) tasks with respective temporal resolutions of $10$ minutes and $1$ hour. In addition, we have formed $3$ disease rate modelling tasks, focusing on the prediction of influenza-like illness (ILI) rates in England (ILI-ENG) and US Health & Human Services (HHS) Regions 2 and 9 (ILI-US2 and ILI-US9). For the ILI rate forecasting tasks, we also introduce the frequency time series of web searches as exogenous predictors. In the ETT and weather tasks, we consider oil and wet-bulb temperature, respectively, as the target variables; we refer to the rest indicators as exogenous variables. We note that the ILI forecasting tasks have considerably more exogenous variables compared to the ETTh1, ETTh2, and weather tasks. Based on the evaluation settings in prior work [32, 28], for the ETT and weather tasks, we set the forecasting horizon $H$ to $96$, $192$, $336$, and $720$ time steps, and use a single test fold of consecutive unseen instances. For the ILI forecasting task, we conduct experiments on $4$ consecutive influenza seasons (2015/16 to 2018/19) as separate test sets ($4$ test folds); the forecasting horizons we consider are $H\!=\!\{7,14,21,28\}$ days ahead. Further details, including the evaluation setup are available in Appendix A.
We compare DeformTime to 6 competitive forecasting models that were initially designed for MTS tasks and, to the best of our knowledge, form the current SOTA methods. These are LightTS [54], DLinear [53], Crossformer [55], PatchTST [32], iTransformer [28], and TimeMixer [44]. We also include a naïve persistence model baseline (Appendix B). For the ILI forecasting tasks, we conduct hyperparameter tuning for all models. In the benchmark tasks, with the exception of Crossformer and LightTS, we adopt the settings from their official repositories to reproduce results. Further details about the experiment settings can be found in Appendix C.
Table 1: Forecasting accuracy results across all tasks and methods. $H$ denotes the forecasting horizon time steps. For the ILI forecasting tasks, the table enumerates the average error across the $4$ test seasons. Complete results per season are shown in Appendix D. $\epsilon\leavevmode\nobreak\ \%$ denotes sMAPE (to reduce column width). The best results are in bold font and the second best are underlined.
Models DeformTime PatchTST iTransformer TimeMixer Crossformer LightTS DLinear Persistence
$H$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$
ETTh1 96 0.1941 14.96 0.2017 15.41 0.2052 15.46 0.2112 16.32 0.2126 16.52 0.2215 17.24 0.2599 20.82 0.2371 18.47
192 0.2116 16.08 0.2409 18.29 0.2429 18.13 0.2382 17.91 0.2820 21.63 0.2636 20.55 0.3798 31.78 0.2803 21.46
336 0.2158 16.27 0.2559 19.29 0.2593 19.11 0.2625 19.72 0.2947 22.65 0.2807 22.15 0.6328 58.34 0.3028 22.90
720 0.2862 21.81 0.3087 23.89 0.2886 22.05 0.3055 23.25 0.3350 24.84 0.5334 44.57 0.7563 69.52 0.3222 25.29
ETTh2 96 0.3121 40.07 0.3145 39.25 0.3420 42.41 0.3454 41.27 0.3486 40.71 0.3507 41.80 0.3349 41.68 0.3522 43.85
192 0.3281 37.90 0.3839 45.45 0.4233 47.44 0.4183 47.49 0.4035 43.16 0.4022 48.01 0.4084 50.67 0.4416 50.24
336 0.3450 37.00 0.4018 46.77 0.4332 45.95 0.4380 46.79 0.4487 49.44 0.4425 51.35 0.4710 55.53 0.4836 53.70
720 0.3640 34.99 0.4960 55.27 0.4565 45.40 0.4729 46.37 0.5832 61.45 0.6252 70.50 0.7981 94.67 0.5199 58.75
Weather 96 0.0244 37.89 0.0258 39.37 0.0277 42.39 0.0322 45.90 0.0271 44.92 0.0293 48.48 0.0251 39.03 0.0329 51.83
192 0.0260 39.33 0.0279 42.02 0.0277 42.77 0.0347 48.62 0.0308 54.14 0.0319 51.45 0.0270 42.68 0.0361 54.92
336 0.0291 44.26 0.0303 45.31 0.0308 46.01 0.0359 49.75 0.0345 62.53 0.0317 50.83 0.0305 47.68 0.0361 55.14
720 0.0363 53.72 0.0389 56.04 0.0395 57.01 0.0457 59.82 0.0395 65.47 0.0386 62.96 0.0352 54.54 0.0394 56.04
ILI-ENG 7 1.6417 28.61 2.3115 27.61 2.3084 26.38 2.1748 25.68 1.8698 25.71 2.2397 52.25 2.8214 43.02 2.1710 24.96
14 2.2308 33.98 3.2547 37.76 3.2301 36.67 3.0209 35.39 2.6543 30.97 2.6879 38.29 3.7922 55.28 3.0625 33.77
21 2.6500 32.70 4.3192 51.11 4.2347 48.93 3.5501 49.36 3.0014 40.57 3.3616 51.78 4.4739 61.25 3.8617 42.03
28 2.7228 40.44 4.9964 59.60 4.8125 55.35 4.1188 54.60 3.1983 46.14 3.4132 55.59 5.0347 67.75 4.5857 49.49
ILI-US2 7 0.4122 16.01 0.7097 24.52 0.6507 23.24 0.5284 20.07 0.4400 16.46 0.4632 16.74 0.7355 27.94 0.6474 22.48
14 0.4752 17.73 0.8635 30.11 0.7896 28.17 0.6556 24.61 0.5852 20.98 0.5827 23.11 0.8435 32.22 0.8135 28.24
21 0.5425 22.13 1.0286 36.70 0.8042 30.03 0.6794 27.68 0.6245 22.29 0.6683 29.27 0.9124 34.93 0.9635 33.51
28 0.5538 22.25 1.1525 42.61 0.9619 36.75 0.8853 36.53 0.6512 23.91 0.7175 27.73 0.9805 37.62 1.1007 38.54
ILI-US9 7 0.2622 12.26 0.4116 19.34 0.4057 18.57 0.3239 15.21 0.3149 14.44 0.3185 15.65 0.4675 23.47 0.4057 18.49
14 0.3084 13.80 0.5020 24.09 0.4702 22.44 0.4060 19.08 0.3571 17.23 0.3791 19.04 0.5467 27.35 0.5008 23.07
21 0.3179 14.24 0.5935 29.40 0.5106 24.11 0.4576 21.40 0.3418 15.90 0.4754 23.74 0.6001 29.66 0.5906 27.41
28 0.3532 15.74 0.6665 33.35 0.6498 31.04 0.5124 24.11 0.3747 16.44 0.4769 23.22 0.6564 32.16 0.6799 31.67
4.2 Forecasting accuracy
Forecasting accuracy for all tasks, models, and horizons is enumerated in Table 1. Mean Absolute Error (MAE) and symmetric Mean Absolute Percentage of Error (sMAPE666sMAPE (Appendix D.1) can provide balanced insights when metrics are averaged across different tasks. or $\epsilon\leavevmode\nobreak\ \%$) are used as our evaluation metrics. Note that for the ILI rate prediction tasks, the error metrics represent the average MAE and sMAPE across $4$ consecutive test seasons (per season results are available in Tables D1D2, and D3, including linear correlation with the target variable as an additional metric), whereas for the remaining benchmark tasks performance metrics are obtained from one train / test split for each forecasting horizon, following previously reported evaluations.
DeformTime displays an overall superior accuracy. It outperforms the most competitive method (that might be a different one per task and horizon), reducing its MAE by $10\%$ and its sMAPE by $7.4\%$ on average across all tasks and forecasting horizons. In more absolute terms, DeformTime is the best performing model in all but one occasion (where it is the second best) based on MAE, and $21$ out of $24$ times based on sMAPE. With the exception of the weather task that draws upon a $1$-year data set from a single location and hence is only expected to offer limited insights while attempting to predict $720$ time steps ($7.5$ days) ahead (see also Appendix A.5), our method offers $>\!10\%$ of MAE reduction in the most challenging forecasting tasks. Specifically, MAE is reduced by $10.5\%$ in the ETT tasks (for $H\!=\!720$ hours), and by $14.9\%$ (England) or $10.3\%$ (US Regions) in the ILI tasks (for $H\!=\!28$ days). Therefore, performance gains do not decrease as the forecasting horizon increases.
When comparing DeformTime to a specific forecaster, we notice that MAE reduction is $>$$15\%$ on average across all tasks and forecasting horizons, ranging from a $15.2\%$ reduction vs. Crossformer to $34.8\%$ vs. DLinear. [^7] A brief note to further explain DLinear’s performance is provided in Appendix D.6. These significant performance gains highlight our model’s capacity to consistently perform well under different tasks and horizons compared to other SOTA models. Interestingly, DLinear fails to surpass the average accuracy of a naïve persistence model (DeformTime reduces persistence’s MAE by $30.1\%$).
Turning our focus to the more interpretable task of predicting ILI rates, we first notice that forecasters that model inter-variable dependencies (Crossformer, LightTS, TimeMixer, iTransformer, and DeformTime) showcase superior performance. Compared to DeformTime’s average sMAPE (across all locations) ranging from $18.96\%$ to $26.14\%$ for $H\!=\!7$ to 2828\!2828 days ahead forecasting respectively, the only consistent competitor is Crossformer ($18.87\%$ to $28.83\%$). The rest of the methods do not perform well as the task becomes more challenging, reaching average sMAPEs ranging from $35.51\%$ to $45.84\%$ for $H\!=\!28$ days. Pending a more detailed evaluation (out-of-scope for this work), there is at least partial evidence to support that DeformTime is a SOTA forecaster for ILI [36, 30, 33, 31]. It not only demonstrates a great regression fit, but also captures the overall trend while forecasting $28$ days ahead (average correlation is >.$>\!.90$, see Tables D1D2, and D3).
We depict the ILI rate forecasts of all models in Appendix G. A snapshot is presented in Figure 2, showcasing $28$ days ahead predictions for influenza season 2018/19 in England. We observe that many models resemble the naïve persistence model, providing smooth but shifted and hence uninformative forecasts. DLinear provides the overall worst fit. With the exception of the iTransformer (see Appendix B for a brief justification), competitive baseline models that capture inter-variable dependencies do perform better. However, their estimates are noisy, and either not capable of capturing the onset (TimeMixer) or intensity (Crossformer and less so LightTS) of the influenza season. Contrary to that, DeformTime provides smoother and more accurate forecasts that corroborate our choice to account for both inter- and intra-variable dependencies.
Figure 2: 28 days ahead forecasting results for influenza season 2018/19 in England (ILI-ENG) for all models. The black line denotes the ground truth, i.e. the reported (actual) ILI rates.
Further experiments, presented in Appendix D.3, provide evidence that DeformTime is also robust to random seed initialisation (Table D4).
4.3 Ablation study for DeformTime
We perform an ablation study to understand the contribution of various operations in DeformTime, namely the V-DAB and T-DAB modules, the position embeddings used in DABs (we conventionally denote both by $\mathbf{P}_{\texttt{v},\texttt{t}}$), and the NAE module with and without position embedding $\mathbf{P}_{\texttt{n}}$. Experiments are conducted on the ILI rate forecasting task for $2$ locations, England (ILI-ENG) and US Region 9 (ILI-US9), and across the $4$ test seasons. Hyperparameters are re-tuned using the same validation approach separately for each ablation variant.
Table 2 enumerates the ablation outcomes, showing the average MAE across all test periods. Evidently, each component or operation contributes to the reduction of MAE. On average, the V-DAB module provides stronger performance improvements ($8.4\%$) compared to the T-DAB module ($4.8\%$). However, for the longest forecasting horizon ($H\!=\!28$) both modules have a similar level of contribution (respectively, $11.4\%$ and $9.8\%$). Hence, establishing inter-variable dependencies is always useful, but intra-variable dependencies become more important in longer forecasting horizons.
Table 2: Ablation study for DeformTime (shortened to DTime). We report the average MAE over the $4$ test periods for tasks ILI-ENG and ILI-US9 across all forecasting horizons ($H$). ‘$\neg$’ denotes the absence of a module or an operation.
$H$ DTime $\neg$ V-DAB $\neg$ T-DAB $\neg$ $\mathbf{P}_{\texttt{v},\texttt{t}}$$\neg$ NAE $\neg$ $\mathbf{P}_{\texttt{n}}$
ILI-ENG 7 1.6417 1.8536 1.8043 1.9420 1.8238 1.8849
14 2.2308 2.4978 2.3570 2.2646 2.4280 2.8982
21 2.6500 2.9173 2.6840 2.8838 2.9523 2.9971
28 2.7228 3.3685 3.2848 3.1082 3.1016 2.8446
Avg. 2.3113 2.6593 2.5325 2.5497 2.5764 2.6562
$H$ DTime $\neg$ V-DAB $\neg$ T-DAB $\neg$ $\mathbf{P}_{\texttt{v},\texttt{t}}$$\neg$ NAE $\neg$ $\mathbf{P}_{\texttt{n}}$
ILI-US9 7 0.2622 0.2709 0.2649 0.2753 0.2665 0.2737
14 0.3084 0.3138 0.3103 0.3174 0.3161 0.3270
21 0.3179 0.3493 0.3236 0.3452 0.3273 0.3355
28 0.3532 0.3664 0.3623 0.3689 0.3638 0.3693
Avg. 0.3104 0.3251 0.3153 0.3267 0.3184 0.3264
The inclusion of the NAE component is equally important as it improves MAE by $6.3\%$ on average. Interestingly, the level of relative improvement increases as the forecasting horizon extends (from $5.8\%$ to $7.6\%$). This potentially highlights that longer forecasting horizons require more abstraction over the feature space, favouring learning from correlated groups of variables. Finally, the use of position embeddings, albeit not entailing a very sophisticated operation, significantly enhances the impact of the V-DAB/T-DAB and NAE modules by $7.2\%$ and $8.9\%$, respectively. We argue that position embeddings, obtained before ($\mathbf{P}_{\texttt{n}}$) and after ($\mathbf{P}_{\texttt{v},\texttt{t}}$) deformation, work in tandem to maintain key information that evidently further improves the overall forecasting accuracy.
4.4 Computational complexity and efficiency of DeformTime
The computational complexity of DeformTime is $\mathcal{O}(L^{2}d+Ld^{2})$, where $L$ denotes the look-back window and $d$ is the size of the hidden layers throughout the method (see Appendix D.5 for a complete derivation). We note that the number of operations for an encoder layer reduces quadratically as we increase $r$ (T-DAB’s time window) to values greater than $1$. Likewise, segmenting the $L$ input time steps to patches of a smaller length (in V-DAB) results in quadratic computational benefits.
Overall, DeformTime has a relatively small (GPU VRAM) memory footprint (see Figure D2). Compared to other transformer-based MTS forecasting methods (Crossformer, PatchTST, and iTransformer), our method exhibits consistently low memory consumption as the number of input variables increases (vs. an exponential increase for the other methods). Furthermore, increasing the look-back window does not overly affect memory consumption either (vs. a more accelerated linear increase for Crossformer and PatchTST). Hence, we deduce that DeformTime is adept to MTS forecasting tasks with a considerable amount of exogenous predictors and longer look-back windows.
5 Conclusions
We propose DeformTime, a novel deep learning architecture for multivariate time series forecasting that uses deformable attention blocks to effectively learn from exogenous predictors, deploying specific operations that aim to capture inter- and intra-variable dependencies. We assess forecasting accuracy using $3$ established benchmark tasks as well as $3$ influenza-like illness prediction tasks with web search frequency time series as exogenous variables, covering $3$ locations (in $2$ countries) and spanning a time period longer than $12$ years. DeformTime yields strong forecasting accuracy across the board, reducing mean absolute error on average by $10\%$ compared to the best baseline model (different for each task and horizon) and by at least $15.2\%$ (for Crossformer [55]) when compared to a specific model. Importantly, performance gains remain stable for longer forecasting horizons. Our experiments, including the ablation study, highlight that modelling variable dependencies is an important attribute. Specifically in the disease forecasting task, where more exogenous predictors are present, the most competitive baselines capture variable dependencies to some extent, whereas models that do not on many occasions cannot surpass the performance of a naïve persistence model (see also Appendix D.7). In contrast to other methods, DeformTime’s GPU memory footprint is not significantly affected by the amount of exogenous variables. We discuss the limitations, societal impact, and future directions for this work in Appendices E and F.
Acknowledgements
V.L. acknowledges all levels of support from the EPSRC grant EP/X031276/1 and a Google donation.
Supplementary Materials
\@bottomtitlebar
Appendix A Data sets for time series forecasting
This section offers a detailed description of the data sets used in our experiments to assess the performance of a series of methods for MTS forecasting. We provide links to data sets that are publicly available. Data sets that were formed by us will be provided in our source code repository. If we cannot share the exact data set due to copyright restrictions, we will provide instructions as to how research teams can obtain access directly from the data provider.
A.1 Established MTS forecasting benchmarks
ETTh1 and ETTh2.
ETTh1 and ETTh2 are electricity transformer temperature data sets that were obtained from $2$ different counties in China [57]. [^8] ETTh1 and ETTh2, github.com/zhouhaoyi/ETDataset Each data set contains $6$ exogenous variables capturing power load attributes, and the target variable which is the temperature of oil. The data set covers a period from July 1, 2016 to June 26, 2018 (although not all data points are used in our experiments to match the setup in related methods that we compare to). The temporal resolution of these data sets is hourly. Adopting the setup in prior work [46, 32, 28], we use a total amount of $14{,}400$ time steps (starting from July 1, 2016) where the first $8{,}640$ are used for training, the next $2{,}880$ for validation, and the last $2{,}880$ for testing.
Weather.
The weather data set [45] contains meteorological measurements collected from a weather station at the Max Planck Biogeochemistry Institute. It covers a year from January 1, 2020 to January 1, 2021.999The data set can be obtained from Autoformer’s repository, github.com/thuml/Autoformer The temporal resolution is $10$ minutes and $20$ meteorological indicators (exogenous variables) are being reported; the target variable is wet-bulb temperature. This results in a total of $52{,}696$ time steps (samples). Based on prior work [55], we use $70\%$, $10\%$, and $20\%$ of the time steps for training, validation, and testing, respectively.
A.2 Forecasting influenza-like illness rates using web search activity
Important components of our data set are shared in the paper’s online repository.1 The shared information allows replication of our work for teams who have obtained the same data-sharing agreements (mainly access to the Google Health Trends API). However, we anticipate that we will be permitted to share all data sets (to make replication and comparisons easier).
Influenza-like illness (ILI) rates.
ILI is defined as the presence of common influenza symptoms (e.g. sore throat, cough, headache, runny nose) in conjunction with high fever. We obtain ILI rates for England (in the United Kingdom) and two US HHS Regions, specifically Region 2 (states of New Jersey and New York) and Region 9 (states of Arizona and California). For England, we obtain ILI rates from the Royal College of General Practitioners (RCGP) that monitors ILI prevalence via an established sentinel network of GP practices throughout the country. An RCGP ILI rate represents the amount of ILI infections in every $100{,}000$ people in the population of England. For the US Regions, data is obtained from the Centers for Disease Control and Prevention (CDC). [^10] US ILI rates (CDC), gis.cdc.gov/grasp/fluview/fluportaldashboard.html An ILI rate from CDC represents the proportion of ILI-related doctor consultations over the total amount of consultations (for any health issue). Hence, the units of the ILI rate in these two monitoring systems (RCGP and CDC) are different. The time span for the obtained ILI rates is the same as the time span for the web search data (see next paragraphs).
Influenza season definition.
An annual influenza season for the HHS US Regions is assumed to start on August 1 of year $\chi$ and end on July 31 of year $\chi\!+\!1$. For England, this is shifted by a month, i.e. an annual influenza season starts on September 1 of year $\chi$ and ends on August 31 of year $\chi\!+\!1$.
Linear interpolation of weekly ILI rates.
ILI rates are reported on a weekly basis. To generate daily data (that increases the amount of training samples by a factor of $7$), we use linear interpolation. It should be noted that for RCGP, Monday determines the start of a week, while for CDC this is determined by Sunday (different ISO specification). We assume that the weekly reported ILI rate is representative of the middle day of a week (Thursday for England and Wednesday for the US Regions) and the ILI rates are then linearly interpolated accordingly.
Real-time delay in ILI rate availability.
We also note that in practice, the reported ILI rate is delayed, i.e. assuming that it is published by RCGP or CDC at time $t$, it actually refers to an ILI rate representing a previous time $t\!-\!\delta$. Specifically, we consider that there is a delay of $\delta\!=\!7$ days for RCGP/England, and a delay of $\delta\!=\!14$ days for CDC/US Regions [42, 36]. This delay, $\delta$, has an impact on the autoregressive time series (of the target variable, in this case, the ILI rate) in a forecasting task (see also section 2). As the delay increases, the forecaster is expected to rely more on the exogenous predictors (that are not delayed) rather than the autoregressive time series of the target variable.
Web search frequency time series.
Web search activity trends, if modelled appropriately [10, 22, 20, 18], are a good indicator of influenza rates in a population and can become a strong exogenous predictor for influenza forecasting [9, 31]. We obtain web search activity data from the Google Health Trends API; this is not a publicly available API, but access can be provided via an application process. [^11] Google Health Trends API access application, support.google.com/trends/contact/trends_api For a day and a certain location, the frequency of a search query is determined as the ratio of searches conducted for a particular term or set of terms divided by the overall search volume. We use a pool of $22{,}071$ unique health-related search queries and obtain their daily frequency from August 1 (US regions) or September 1 (England), 2006 to July 31 (US regions) or August 31 (England), 2019. However, we note that not all search queries are used as exogenous variables in our forecasting models. A feature selection process is described in Appendix A.3.
Given that for the US, we can only obtain data at the state level (as opposed to regional), we use a weighted average of the state-level search query frequencies. The weights are based on the population of each state. In particular, for Region 2, the states of New Jersey and New York have weights of $0.32$ and $0.68$, whereas for Region 9, the states of Arizona and California have weights of $0.16$ and $0.84$, respectively. Note that smaller locations (or distant ones) that may be part of an HHS Region are excluded from the web search data sets; given their relatively small population, these locations do not have a significant impact on the reported ILI rates and the data obtained from the Google Health Trends API are quite sparse.
Training, validation, and test sets.
We assess the accuracy of forecasters across the last $4$ influenza seasons ($4$ separate test sets) in England and the US regions. Each test period is a complete influenza season (as previously defined). For each test period, we train models based on the $9$ influenza seasons that precede it. A part of each training set is used for validating model decisions, including hyperparameters. We construct validation sets using the following strategy. Each validation set has $180$ days in total, using $60$-day periods to capture the onset, peak, and the period after the peak (or the outset) of an influenza season. Each $60$-day period comes from a different influenza season in the training set; we use the last $3$ influenza seasons in the training data to make sure our validation process has a recency effect. The last, penultimate, and third from last influenza seasons are used to define the outset, peak, and onset validation periods, respectively. We use CDC’s definition to determine the onset of an influenza season, i.e. a time point is deemed to be the onset when the subsequent ILI rates exceed a threshold for $2$ consecutive weeks. [^12] CDC’s influenza season onset definition, cdc.gov/flu/weekly/overview.htm The peak point is simply the highest ILI rate within an influenza season. The outset has the inverse definition from the onset, i.e. the last time point where the ILI rate exceeds the onset threshold for two consecutive weeks. Once these time points are determined, they become the $30$th point in a $60$-day validation period. Figure A1 shows an example of the validation periods that were determined when training an ILI rate model for England using (training) data from September 2009 to August 2018 (with the test period being September 2018 to August 2019).
Figure A1: An example of how the training, validation, and test sets are constructed when the test influenza season is 2018/19 (England). The lines in blue, red, and orange colour denote the training, validation, and test periods, respectively. To form the validation set from our training data, we select the period after the peak (outset) from the third to last influenza season, the period around the peak from the penultimate season, and a period around influenza onset from the last season.
A.3 Feature selection for web search activity time series
Originally, we consider $22{,}071$ search queries. We then perform two feature selection steps to maintain more relevant queries to the ILI forecasting task. First, we apply a semantic filter to remove queries that are not related to the topic of influenza, similarly to the approach presented by Lampos et al. [21]. We obtain an embedding representation for each search query using a pre-trained sentence BERT model. [^13] Sentence BERT, huggingface.co/bert-base-uncased We also obtain the embeddings of influenza-related (base) terms and expressions, such as “flu”, “flu symptoms” etc. We compute the cosine similarity between each base term and search query and maintain the top-$1000$ (England) or top-$500$ (US regions) search queries (per base term). This leaves $4{,}396$ and $2{,}479$ queries respectively for England and the US regions. [^14] Base and filtered search terms are provided in DeformTime’s repository (see footnote 1). The second selection step is dynamic, i.e. it might have a different outcome as the training data change. In this step, we compute the linear correlation between the remaining search queries and the target variable (ILI rate) in the $5$ most recent influenza seasons in the training data (to impose recency). We maintain search queries with a correlation that exceeds a correlation threshold $\tau$, which is a learnable hyperparameter (see also Appendix C).
A.4 Data normalisation
In all forecasting tasks, we standardise all variables, an operation also known as z-scoring (zero mean, unit standard deviation), each time based on the training data. The prediction output is de-normalised back to its original scale prior to being compared with the (de-normalised) ground truth. In addition to that, for the ETT and weather data sets, we also z-score each variable within the input’s look-back window, in accordance with previous work [27, 58, 32].
A.5 Criticism of data sets used to benchmark forecasting methods
Although we present results using the ETTh1, ETTh2, and weather data sets to compare with other forecasting methods in the literature, we consider these data sets to have several shortcomings. Clearly, a weather data set that covers only a $1$-year period and is based on the sensors from just one geographical location cannot be expected to provide good enough insights into meteorological forecasting. It captures a negligible portion of meteorological data and, as a result, derived forecasting models cannot have any practical impact (compared to actual models that are used in meteorology). Hence, using this data set for training and evaluating a forecasting model may, unfortunately, result in misleading or at least inconclusive outcomes. Similar issues, but perhaps to a smaller extent, are present in the ETTh1 and ETTh2 data sets. On the one hand, these data sets seem to have a more solid practical application. However, experiments on their ETTm1 and ETTm2 variants, where the temporal resolution changes from $1$ hour to $15$ minutes (same data, shorter temporal resolution), showcase part of the problem. When we compared the forecasting performance of SOTA baselines between the two different temporal resolutions, we surprisingly found that the prediction based on the hourly sampled data set has better accuracy compared to the quarter-hourly sampled data set even at longer forecasting horizons. For example, models that conduct $192$ or even $336$ time steps (equivalent to $8$ or $14$ days) ahead forecasting with ETTh1 and ETTh2 achieve a lower MAE compared to a $720$ time steps ahead forecast with ETTm1 and ETTm2 (equivalent to $7.5$ days ahead forecast), respectively [32, 46, 29].
We also noticed that some papers run experiments on US national weekly ILI rates obtained from the CDC [59, 43, 32]. The first concerning observation was that in these experiments, the input variables were based on variates (columns) from the CDC extract (spreadsheet) that only provide redundant insights compared to the ILI rate (e.g. the raw numerator and denominator of an ILI rate or the population unweighted ILI rate). In addition, some forecasting horizons that were explored in these benchmarks were practically unreasonable (even from a biological perspective, there are limits in predicting the future, especially when it comes to viruses). SOTA forecasting models for influenza generally provide good accuracy at approximately a $2$ weeks ahead forecasting horizon [33, 31]. Anything beyond that with satisfactory performance should be considered as a very important development within the disease modelling community. We do think DeformTime might fall in that category as it delivers good results for $3$ or $4$ weeks ahead forecasting horizons. Contrary to that, in the aforementioned ILI rate benchmark, there were forecasting horizons of $24$ or $60$ weeks ahead (i.e. even more than a year!). Of course, the accuracy of the forecasting models under these forecasting horizons was reportedly detrimentally poor (influenza does not have a strong periodicity), and consequently could not be used for any meaningful comparative conclusions. If all forecasting models are performing poorly, even the least poorly performing one still is an inadequate forecaster (for the underlying task).
Beyond this, we have also excluded data sets, such as traffic [45] and electricity consumption [57], that do not have one explicitly defined target variable, but instead focus on multivariate predictions (multi-task learning). Consequently, these data sets are not entirely compatible with our forecasting task or method (DeformTime makes predictions about one target variable).
To compensate for that we have developed a set of $3$ influenza forecasting tasks. Models from these tasks can find direct real-world applications, i.e. become part of influenza monitoring systems in England or the US. The evaluation process is rigorous, i.e. across $4$ consecutive influenza seasons as opposed to using $1$ fixed test period that may again lead to biased insights.
Appendix B Summary of other forecasting models
Here we provide a list of the models that we are comparing DeformTime to:
  • LightTS [54] uses MLP layers as the building blocks to extract both inter-variable and intra-variable dependencies upon time series with different temporal resolutions.
  • DLinear [53] uses a fully connected layer along the temporal dimension with seasonal-trend decomposition to regress the historical values for future predictions.
  • Crossformer [55] is a transformer-based model that divides the input into patches and proposes a two-stage attention layer to capture inter- and intra-variable dependencies.
  • PatchTST [32] is a transformer-based model that takes segmented patches as input tokens for the model. The prediction of each variable is designed to be independent of one another, i.e. with no inter-variable dependencies established.
  • iTransformer [28] is a transformer-based model that embeds the sequence dimension rather than the input variable dimension. It currently provides the SOTA performance on the long-term forecast tasks. [^15] Long-term forecast tasks ranking, github.com/thuml/Time-Series-Library We note that although iTransformer captures inter-variable dependencies, it achieves that by embedding the input along the sequence dimension. This disrupts the temporal structure within each variable, leading to a sub-optimal performance in our empirical evaluation.
  • TimeMixer [44] is an MLP-based model that predicts the seasonal and trend components at different sampling scales and mixes forecasts to form the final prediction. Note that the authors consider variable mixing optional for their model. We opted to deploy the version that conducts variable mixing for the ILI rate forecasting tasks.
  • The persistence model is a naïve baseline that uses the last seen value of the target variable in the input as the forecast.
Appendix C Supplementary experiment settings
This section is supplementary to section 4.1 in the main paper. All experiments were conducted (in parallel) using a server with 2 NVIDIA A40 GPUs.
C.1 Input variable re-arrangement based on correlation with the target variable
In the ETT and weather forecasting tasks, we rearrange variables (see NAE in section 3.2) based on their linear correlation with the target variable across all time steps preceding the validation set. In the ILI rate forecasting tasks where training sets cover a period of $9$ years prior to the test period, we only use the last $5$ influenza seasons in the training set to obtain correlations. This is because both user search behaviour and search engine characteristics (e.g. automatic search suggestions) evolve, and this approach to feature selection reinforces a weak recency effect that is ultimately beneficial to prediction accuracy [50, 31].
C.2 DeformTime setup
We use $G=4$ variable neighbourhoods throughout (see section 3.2). This setting ($G=4$) also determines the number of group partitions for offset generation and multi-head attention in T-DAB (section 3.4). The learnable amplitude $\alpha\!\in\!\{3,5,7,9\}$ (see sections 3.3 and 3.4) is tuned collectively for both V-DAB and T-DAB. We keep the kernel size $k$ of the interpolation functions $\phi(\cdot)$ and $\psi(\cdot)$ identical to $\alpha$ throughout the experiments. In all experiments, DeformTime has $2$ encoder layers (see section 3.5). For the ETT and weather tasks, we set $r\!=\!1$ in the first encoder layer and select from $r\!\in\!\{6,12,24\}$ time steps in the second encoder layer (see section 3.4). In addition, we select the segmentation length from $\ell\!\in\!\{6,12,24\}$ time steps (see section 3.3) collectively for both encoder layers. Given the more direct weekly temporal resolution in the reported ILI rates and the fact that web search activity trends do have weekly patterns, in the ILI forecasting task, we set $\ell\!=\!7$ time steps (days), and also set $r\!=\!1$ and $r\!=\!7$ time steps in the first and second encoder layer, respectively. The segmentation stride $s$ is kept the same as $\ell$ throughout the experiments.
C.3 Optimisation and hyperparameter settings
The look-back window $L$ is set to $336$ time steps for the ETTh1, ETTh2, and weather forecasting tasks, based on Nie et al. [32]. For influenza forecasting, we set $L$ to $\{28,28,56,56\}$ days for forecasting horizons $H\!=\!\{7,14,21,28\}$ days ahead, respectively. Only for influenza forecasting, we also set the batch size and $d$ to $64$. For the other tasks, these become learnable parameters (see next section). The number of training epochs is set to $100$ for ETTh1, ETTh2, and weather tasks and $50$ for the influenza forecasting task. Neural networks are optimised with Adam using mean squared error loss on all outputs. Note that the output can be a time series (sequence) of one variable (the target) or more variables (target and exogenous) depending on the forecasting method. As an exception, we use MAE loss for all models tuned on the weather forecasting task. Training stops early when the validation error of the target variable at time $t\!+\!H$ does not decrease further (compared to its lowest value) for $5$ consecutive epochs. We then use the model with the lowest validation loss.
For the ILI rate forecasting tasks, all models are trained with a fixed random seed equal to ‘42’. For ETTh1, ETTh2, and weather tasks, we fix the seed to ‘2021’ based on previous work [32, 53, 44], with the exception of iTransformer where we set the seed to ‘2023’ in accordance with the official configuration.
C.4 Hyperparameter validation
[…]
Appendix D Supplementary results
Table D1: ILI rate forecasting accuracy results in England (ILI-ENG) using web search frequency time series as exogenous variables. $\rho$ and $\epsilon\leavevmode\nobreak\ \%$ denote linear correlation and sMAPE, respectively. The best results are in bold font and the second best are underlined.
2015/16 2016/17 2017/18 2018/19 Average
$H$ Model $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$
7 Persistence 0.9072 1.7077 22.43 0.9129 1.2287 22.67 0.8539 4.0441 31.08 0.8944 1.7037 23.67 0.8921 2.1710 24.96
DLinear 0.8710 2.6985 38.86 0.8769 1.6879 33.77 0.8231 4.2734 50.14 0.8906 2.6258 49.30 0.8654 2.8214 43.02
PatchTST 0.8832 1.9823 26.13 0.8744 1.4307 25.39 0.8475 3.7925 31.23 0.8526 2.0403 27.70 0.8644 2.3115 27.61
iTransformer 0.8773 2.0645 27.26 0.8911 1.3734 24.06 0.8180 4.1077 31.26 0.8911 1.6881 22.94 0.8694 2.3084 26.38
TimeMixer 0.9056 1.7179 22.27 0.8616 1.4676 24.97 0.8694 3.4828 28.09 0.8388 2.0310 27.39 0.8688 2.1748 25.68
LightTS 0.8385 2.2625 38.87 0.9432 2.4880 89.24 0.9441 2.6512 52.01 0.9426 1.5570 28.87 0.9171 2.2397 52.25
Crossformer 0.8879 2.3314 36.08 0.9591 0.8606 17.82 0.9234 2.9133 27.04 0.9436 1.3737 21.89 0.9285 1.8698 25.71
DeformTime 0.9728 1.3786 19.57 0.9665 1.3083 29.49 0.9585 2.6313 41.40 0.9436 1.2485 23.96 0.9603 1.6417 28.61
14 Persistence 0.8292 2.3161 30.16 0.8414 1.6867 29.54 0.7289 5.8180 42.30 0.8031 2.4291 33.08 0.8006 3.0625 33.77
DLinear 0.7810 3.6354 52.29 0.7857 2.1872 42.47 0.6834 5.5368 60.91 0.8351 3.8092 65.47 0.7713 3.7922 55.28
PatchTST 0.8078 2.6267 35.59 0.7730 1.9562 33.55 0.6722 5.6075 43.64 0.7218 2.8284 38.25 0.7437 3.2547 37.76
iTransformer 0.7731 2.7620 36.46 0.7776 1.9929 34.17 0.6576 5.7796 44.34 0.7850 2.3862 31.71 0.7483 3.2301 36.67
TimeMixer 0.7706 2.6783 34.25 0.8684 1.5621 29.06 0.7164 5.0940 39.59 0.7503 2.7491 38.66 0.7764 3.0209 35.39
LightTS 0.8109 2.5494 32.28 0.9233 1.6292 36.46 0.7821 4.4367 39.46 0.8915 2.1365 44.97 0.8519 2.6879 38.29
Crossformer 0.7940 2.9124 41.75 0.8843 1.4711 23.39 0.8470 3.8295 30.88 0.9007 2.4044 27.86 0.8565 2.6543 30.97
DeformTime 0.9259 2.0556 25.36 0.9400 1.3180 33.86 0.8964 3.7631 47.33 0.9154 1.7863 29.36 0.9194 2.2308 33.98
21 Persistence 0.7357 2.8604 36.98 0.7465 2.1552 36.45 0.6002 7.3119 52.46 0.6952 3.1193 42.23 0.6944 3.8617 42.03
DLinear 0.4466 4.1958 56.02 0.6441 2.8294 52.56 0.4920 6.9514 71.66 0.6608 3.9192 64.75 0.5609 4.4739 61.25
PatchTST 0.6126 3.6607 48.57 0.6214 2.6377 44.77 0.4427 7.3828 60.54 0.5861 3.5956 50.55 0.5657 4.3192 51.11
iTransformer 0.5310 4.1390 55.62 0.5987 2.6718 43.65 0.4403 7.0816 55.17 0.6556 3.0464 41.29 0.5564 4.2347 48.93
TimeMixer 0.5398 3.8805 56.71 0.7799 1.7940 30.45 0.7497 5.5338 58.68 0.7539 2.9919 51.58 0.7058 3.5501 49.36
LightTS 0.7082 3.3630 45.16 0.8696 1.9335 58.65 0.8132 5.6422 51.74 0.8594 2.5078 51.59 0.8126 3.3616 51.78
Crossformer 0.8113 2.6710 36.42 0.8014 1.8035 31.29 0.8504 4.5615 49.08 0.7331 2.9697 45.48 0.7991 3.0014 40.57
DeformTime 0.8819 2.1700 27.01 0.8859 1.8980 32.65 0.9110 4.8984 44.78 0.9092 1.6335 26.36 0.8970 2.6500 32.70
28 Persistence 0.6408 3.3786 43.17 0.6344 2.6018 42.68 0.4813 8.5576 60.33 0.5733 3.8047 51.79 0.5825 4.5857 49.49
DLinear 0.3917 4.5039 60.30 0.5078 3.3928 60.94 0.3911 7.7738 77.58 0.5774 4.4683 72.19 0.4670 5.0347 67.75
PatchTST 0.4271 4.5240 60.22 0.4477 3.2658 53.01 0.3203 8.0693 65.91 0.4486 4.1265 59.27 0.4109 4.9964 59.60
iTransformer 0.3492 4.9542 65.45 0.6761 2.2745 38.38 0.2947 8.3463 69.39 0.5286 3.6750 48.18 0.4622 4.8125 55.35
TimeMixer 0.5395 3.8587 57.17 0.6890 2.4063 45.76 0.4882 6.7783 62.93 0.6632 3.4318 52.52 0.5950 4.1188 54.60
LightTS 0.5812 3.4254 43.91 0.8311 1.9020 36.26 0.7151 6.0675 83.54 0.8900 2.2578 58.66 0.7543 3.4132 55.59
Crossformer 0.7299 3.1253 41.56 0.8082 1.8197 33.57 0.7841 5.4882 55.97 0.9349 2.3599 53.48 0.8143 3.1983 46.14
DeformTime 0.7919 2.6540 49.29 0.9229 1.8100 41.87 0.8493 4.8324 37.53 0.9499 1.5947 33.06 0.8785 2.7228 40.44
D.1 Symmetric Mean Absolute Percentage of Error
To compare results across different tasks that have different units, we use the sMAPE metric. We note the general (and correct) perception that occasionally sMAPE may mislead (by over- or under-estimating error), and hence it can only be used for partial insights in conjunction with a more robust error metric. Hence, our main error metric is MAE. For the ILI rate forecasting tasks, we additionally show linear correlation as an established metric in related literature [10, 20, 50].
We use the following definition of sMAPE. For a series of estimates $\hat{\mathbf{y}}\!\in\!\mathbb{R}^{n}$ and a corresponding series of true values $\mathbf{y}\!\in\!\mathbb{R}^{n}$, sMAPE is given by
$\texttt{sMAPE}\!\left(\hat{\mathbf{y}},\mathbf{y}\right)=\frac{100}{n}\sum_{j=1}^{n}\frac{\left|\hat{y}_{j}-y_{j}\right|}{0.5\left(\left|\hat{y}_{j}\right|+\left|y_{j}\right|\right)}\,.$ (12)
In Tables throughout the manuscript, we denote sMAPE by using $\epsilon\leavevmode\nobreak\ \%$.
Table D2: ILI rate forecasting accuracy results in US Region 2 (ILI-US2) using web search frequency time series as exogenous variables. $\rho$ and $\epsilon\leavevmode\nobreak\ \%$ denote linear correlation and sMAPE, respectively. The best results are in bold font and the second best are underlined.
2015/16 2016/17 2017/18 2018/19 Average
$H$ Model $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$
7 Persistence 0.7758 0.4114 22.26 0.8271 0.6803 24.57 0.7589 1.0284 25.48 0.8954 0.4696 17.59 0.8143 0.6474 22.48
DLinear 0.7456 0.4982 29.23 0.7980 0.8210 31.85 0.7133 1.0910 30.82 0.8955 0.5318 19.84 0.7881 0.7355 27.94
PatchTST 0.7666 0.4290 23.78 0.7973 0.7546 26.97 0.6937 1.1018 26.52 0.8624 0.5533 20.79 0.7800 0.7097 24.52
iTransformer 0.7986 0.3865 21.30 0.8171 0.7367 27.07 0.7784 0.9080 23.22 0.8545 0.5717 21.38 0.8122 0.6507 23.24
TimeMixer 0.8237 0.3842 21.76 0.8710 0.6449 24.16 0.9124 0.6255 17.87 0.8877 0.4588 16.49 0.8737 0.5284 20.07
LightTS 0.7930 0.4045 23.23 0.9056 0.5806 18.01 0.9359 0.5688 15.39 0.9451 0.2989 10.33 0.8949 0.4632 16.74
Crossformer 0.8796 0.3382 21.39 0.9107 0.4975 18.32 0.9009 0.6309 15.61 0.9523 0.2936 10.52 0.9109 0.4400 16.46
DeformTime 0.8887 0.3428 21.86 0.9463 0.4796 18.66 0.9008 0.5369 12.88 0.9622 0.2894 10.64 0.9245 0.4122 16.01
14 Persistence 0.6872 0.5000 26.67 0.7617 0.8545 31.23 0.6331 1.3027 32.42 0.8436 0.5966 22.65 0.7314 0.8135 28.24
DLinear 0.6546 0.5761 33.62 0.7335 0.9406 36.90 0.6007 1.2213 34.62 0.8261 0.6360 23.74 0.7037 0.8435 32.22
PatchTST 0.6628 0.5103 27.27 0.7350 0.9267 34.44 0.5993 1.3085 31.82 0.7922 0.7085 26.90 0.6973 0.8635 30.11
iTransformer 0.6864 0.4894 26.35 0.7548 0.8934 33.26 0.6025 1.1639 30.28 0.8294 0.6115 22.78 0.7183 0.7896 28.17
TimeMixer 0.7741 0.4438 25.01 0.8403 0.7314 27.32 0.8043 0.8911 25.13 0.8622 0.5560 20.99 0.8202 0.6556 24.61
LightTS 0.7203 0.5086 28.62 0.8601 0.7079 22.95 0.9191 0.6698 23.47 0.9409 0.4445 17.40 0.8601 0.5827 23.11
Crossformer 0.7872 0.4357 24.61 0.8640 0.6502 22.95 0.8057 0.8159 20.83 0.9241 0.4389 15.54 0.8453 0.5852 20.98
DeformTime 0.8050 0.4404 23.82 0.9271 0.5029 18.53 0.9126 0.6351 16.34 0.9652 0.3226 12.21 0.9025 0.4752 17.73
21 Persistence 0.5918 0.5680 30.04 0.7017 1.0055 37.20 0.5137 1.5443 38.93 0.7792 0.7361 27.88 0.6466 0.9635 33.51
DLinear 0.6016 0.6023 35.11 0.6897 0.9671 38.06 0.4966 1.3272 38.76 0.7936 0.7530 27.80 0.6454 0.9124 34.93
PatchTST 0.5775 0.5592 29.41 0.6286 1.1661 43.41 0.5189 1.4395 38.13 0.6478 0.9495 35.86 0.5932 1.0286 36.70
iTransformer 0.6418 0.4790 24.54 0.7515 0.8784 34.22 0.8230 1.0845 31.85 0.7321 0.7748 29.52 0.7371 0.8042 30.03
TimeMixer 0.7368 0.4655 27.79 0.8281 0.7853 31.00 0.8509 0.9275 30.07 0.8803 0.5394 21.86 0.8240 0.6794 27.68
LightTS 0.8484 0.5606 32.47 0.8936 0.6749 34.80 0.9177 0.8065 28.47 0.8951 0.6310 21.34 0.8887 0.6683 29.27
Crossformer 0.7729 0.4634 26.54 0.8996 0.6688 25.62 0.7621 0.8882 23.17 0.9354 0.4777 13.82 0.8425 0.6245 22.29
DeformTime 0.7414 0.4568 25.94 0.9028 0.6189 26.44 0.9313 0.6981 21.04 0.9408 0.3963 15.09 0.8791 0.5425 22.13
28 Persistence 0.4860 0.6393 33.84 0.6374 1.1447 42.55 0.4046 1.7516 44.79 0.7071 0.8674 32.98 0.5588 1.1007 38.54
DLinear 0.4836 0.6512 37.96 0.6112 1.0328 41.19 0.4125 1.4045 40.27 0.6840 0.8337 31.08 0.5479 0.9805 37.62
PatchTST 0.4013 0.6678 36.55 0.5274 1.2615 48.85 0.3903 1.6234 44.67 0.5579 1.0576 40.39 0.4692 1.1525 42.61
iTransformer 0.5258 0.5949 32.54 0.6567 1.0632 41.38 0.5726 1.3053 38.82 0.6738 0.8842 34.26 0.6072 0.9619 36.75
TimeMixer 0.5756 0.5835 36.06 0.7251 1.0106 39.21 0.7694 1.2470 43.49 0.7467 0.7000 27.37 0.7042 0.8853 36.53
LightTS 0.8262 0.6018 33.56 0.8332 0.8180 31.54 0.8556 0.8913 27.03 0.9011 0.5589 18.78 0.8540 0.7175 27.73
Crossformer 0.7102 0.4697 27.38 0.9064 0.7624 28.93 0.8432 0.8253 20.69 0.8782 0.5474 18.64 0.8345 0.6512 23.91
DeformTime 0.8369 0.4716 27.21 0.9562 0.6017 25.66 0.9253 0.5954 17.62 0.9294 0.5463 18.52 0.9119 0.5538 22.25
Table D3: ILI rate forecasting accuracy results in US Region 9 (ILI-US9) using web search frequency time series as exogenous variables. $\rho$ and $\epsilon\leavevmode\nobreak\ \%$ denote linear correlation and sMAPE, respectively. The best results are in bold font and the second best are underlined.
2015/16 2016/17 2017/18 2018/19 Average
$H$ Model $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$ $\rho$ MAE $\epsilon\leavevmode\nobreak\ \%$
7 Persistence 0.8384 0.3945 18.58 0.8494 0.2791 16.25 0.7304 0.5949 21.55 0.8898 0.3543 17.57 0.8270 0.4057 18.49
DLinear 0.8071 0.4769 25.07 0.8092 0.3727 22.02 0.6759 0.5704 22.96 0.8925 0.4499 23.82 0.7962 0.4675 23.47
PatchTST 0.8248 0.3993 19.30 0.8318 0.2802 16.91 0.7068 0.6078 23.25 0.8796 0.3592 17.91 0.8107 0.4116 19.34
iTransformer 0.8380 0.3736 17.55 0.8446 0.2721 16.22 0.6664 0.6856 26.00 0.9230 0.2915 14.52 0.8180 0.4057 18.57
TimeMixer 0.9079 0.2892 13.61 0.8521 0.2733 16.29 0.7923 0.4896 18.47 0.9525 0.2435 12.48 0.8762 0.3239 15.21
LightTS 0.8896 0.2790 13.63 0.8858 0.2931 16.95 0.8277 0.4474 16.82 0.9433 0.2543 15.20 0.8866 0.3185 15.65
Crossformer 0.9305 0.2556 12.21 0.9524 0.2874 14.89 0.8213 0.4524 17.08 0.9435 0.2640 13.58 0.9119 0.3149 14.44
DeformTime 0.9161 0.2437 12.46 0.9356 0.2364 11.69 0.8744 0.3664 13.61 0.9675 0.2023 11.29 0.9234 0.2622 12.26
14 Persistence 0.7572 0.4875 23.00 0.7869 0.3385 19.84 0.6439 0.7283 26.96 0.8239 0.4488 22.49 0.7530 0.5008 23.07
DLinear 0.7323 0.5767 29.99 0.7371 0.4384 25.44 0.6105 0.6635 27.69 0.8216 0.5085 26.27 0.7254 0.5467 27.35
PatchTST 0.7626 0.4810 23.33 0.7411 0.3524 21.48 0.6246 0.7164 28.11 0.8102 0.4583 23.44 0.7346 0.5020 24.09
iTransformer 0.7837 0.4310 20.63 0.7629 0.3448 20.49 0.6138 0.6853 27.21 0.8495 0.4196 21.41 0.7525 0.4702 22.44
TimeMixer 0.8319 0.3600 17.41 0.8221 0.3153 18.22 0.6668 0.6111 23.48 0.9139 0.3377 17.21 0.8087 0.4060 19.08
LightTS 0.8697 0.4018 20.65 0.8237 0.3418 20.50 0.7802 0.4999 20.31 0.9366 0.2730 14.70 0.8525 0.3791 19.04
Crossformer 0.8787 0.3333 16.05 0.8965 0.2980 16.74 0.8339 0.4325 16.91 0.9169 0.3647 19.21 0.8815 0.3571 17.23
DeformTime 0.9033 0.2962 14.39 0.9374 0.2893 13.15 0.8490 0.3897 14.62 0.9378 0.2584 13.05 0.9069 0.3084 13.80
21 Persistence 0.6682 0.5831 27.54 0.7150 0.4013 23.26 0.5734 0.8466 31.93 0.7520 0.5313 26.92 0.6772 0.5906 27.41
DLinear 0.6687 0.6032 30.66 0.6920 0.4960 28.49 0.5830 0.7252 30.15 0.7778 0.5760 29.33 0.6804 0.6001 29.66
PatchTST 0.6899 0.5705 28.41 0.6399 0.4255 26.14 0.5214 0.7963 32.80 0.7137 0.5815 30.27 0.6412 0.5935 29.40
iTransformer 0.8659 0.3645 17.75 0.7078 0.3702 21.73 0.4561 0.8642 33.94 0.8298 0.4433 23.01 0.7149 0.5106 24.11
TimeMixer 0.8280 0.4112 19.07 0.8180 0.3058 17.67 0.6555 0.6812 25.45 0.8634 0.4322 23.41 0.7912 0.4576 21.40
LightTS 0.8354 0.4252 20.19 0.7748 0.4021 25.39 0.6920 0.6201 25.57 0.8699 0.4542 23.80 0.7930 0.4754 23.74
Crossformer 0.9565 0.3267 15.49 0.8669 0.3168 15.95 0.8777 0.3964 15.11 0.9135 0.3273 17.07 0.9036 0.3418 15.90
DeformTime 0.8331 0.3543 15.62 0.8966 0.3039 14.99 0.9141 0.3536 13.06 0.9249 0.2598 13.27 0.8922 0.3179 14.24
28 Persistence 0.5763 0.6781 32.15 0.6261 0.4733 27.09 0.5059 0.9541 36.29 0.6705 0.6140 31.14 0.5947 0.6799 31.67
DLinear 0.6083 0.6496 32.81 0.6016 0.5329 30.18 0.5506 0.7893 32.75 0.7438 0.6539 32.89 0.6261 0.6564 32.16
PatchTST 0.6067 0.6457 32.75 0.5192 0.4903 30.18 0.4424 0.8809 36.70 0.6555 0.6492 33.78 0.5560 0.6665 33.35
iTransformer 0.6984 0.5531 27.80 0.5779 0.4444 25.70 0.3211 0.9942 40.08 0.6797 0.6077 30.60 0.5693 0.6498 31.04
TimeMixer 0.8130 0.4056 21.18 0.8184 0.3244 19.25 0.6444 0.7787 31.60 0.7816 0.5409 24.41 0.7644 0.5124 24.11
LightTS 0.8462 0.4765 24.49 0.6952 0.4834 25.81 0.7176 0.5616 22.25 0.9037 0.3862 20.33 0.7907 0.4769 23.22
Crossformer 0.8969 0.3398 16.47 0.8346 0.3647 16.99 0.8514 0.4365 15.42 0.8923 0.3578 16.87 0.8688 0.3747 16.44
DeformTime 0.8888 0.3718 15.55 0.9046 0.3153 14.57 0.9037 0.4185 16.18 0.9260 0.3069 16.68 0.9058 0.3532 15.74
D.2 Detailed ILI forecasting results
Complete results (all test periods, models, and forecasting horizons) in the ILI rate forecasting tasks for England (ILI-ENG), US Region 2 (ILI-US2), and US Region 9 (ILI-US9) are presented in Tables D1D2, and D3, respectively. In addition to MAE and sMAPE, we also show the linear correlation (denoted by $\rho$) between estimates and the target variable throughout each test period.
D.3 Seed robustness for DeformTime
We have examined the robustness of DeformTime across different seeds with hyperparameter tuning by re-running the experiments on the ILI-ENG data set across all forecasting horizons and test sets. The results are enumerated in Table D4. The seed used for the results presented in the main paper (‘42’) provided a performance that did not deviate significantly compared to other seeds. Hence, we conclude that DeformTime is robust to random seed initialisation. We note that compared to the average performance across the explored seeds, seed ‘42’ provides a rather conservative estimate.
Table D4: Seed control ($5$ seeds) for DeformTime using the ILI-ENG data set across all forecasting horizons ($H$). $\mu$ and $\sigma$ denote the mean and standard deviation of the $5$ obtained MAEs per test period. The results in the main paper were obtained for seed ‘42’.
$H$ seed 2015/16 2016/17 2017/18 2018/19 Average
7 42 1.379 1.308 2.631 1.248 1.642
10 1.416 1.063 2.266 1.360 1.526
111 1.386 0.886 2.315 1.690 1.569
$1$ 1.476 1.404 1.986 1.533 1.600
1234 1.443 1.109 2.197 1.226 1.494
$\mu\leavevmode\nobreak\ (\sigma)$ 1.42 (0.04) 1.15 (0.18) 2.28 (0.21) 1.41 (0.18) 1.57 (0.06)
14 42 2.056 1.318 3.763 1.786 2.231
10 1.696 1.514 3.439 1.712 2.090
111 1.861 1.297 3.556 1.842 2.139
$1$ 1.621 1.495 3.327 1.396 1.960
1234 1.837 1.137 3.879 1.391 2.061
$\mu\leavevmode\nobreak\ (\sigma)$ 1.81 (0.15) 1.35 (0.14) 3.59 (0.20) 1.63 (0.19) 2.10 (0.10)
21 42 2.170 1.898 4.898 1.633 2.650
10 2.079 1.507 4.871 1.874 2.583
111 2.354 1.825 4.668 1.944 2.697
$1$ 2.236 1.943 4.158 1.625 2.491
1234 2.147 1.595 4.406 1.842 2.498
$\mu\leavevmode\nobreak\ (\sigma)$ 2.20 (0.09) 1.75 (0.17) 4.60 (0.28) 1.78 (0.13) 2.58 (0.09)
28 42 2.654 1.810 4.832 1.595 2.723
10 2.486 1.692 5.194 1.807 2.795
111 2.655 1.825 4.887 1.907 2.819
$1$ 2.251 1.799 4.529 2.082 2.665
1234 2.692 1.808 4.939 1.613 2.763
$\mu\leavevmode\nobreak\ (\sigma)$ 2.55 (0.17) 1.79 (0.05) 4.88 (0.21) 1.80 (0.18) 2.75 (0.06)
D.4 Ablation study for DeformTime – Additional information
Here we provide additional details to supplement the ablation study presented in section 4.3.
$\neg$ V-DAB: Denotes that we train a DeformTime model without the V-DAB encoder branch (see section 3.3). Within each encoder layer, we only use the T-DAB branch for sequence encoding. Instead of projecting the concatenated output of two encoder branches, we take $\mathbf{Z}_{\texttt{c}}$ directly from T-DAB as the output $\mathbf{Z}_{j}$ of the $j$-th encoder layer.
$\neg$ T-DAB: Denotes that we train a DeformTime model without the T-DAB encoder branch (see section 3.4). We take $\mathbf{Z}_{\texttt{c}}$ directly from V-DAB as the output $\mathbf{Z}_{j}$ of the $j$-th encoder layer.
$\neg\leavevmode\nobreak\ \mathbf{P}_{\texttt{v},\texttt{t}}$: Denotes that we do not use the relative positional biases (both $\mathbf{P}_{\texttt{v}}$ and $\mathbf{P}_{\texttt{t}}$) when obtaining the value embedding for both the V-DAB and T-DAB modules.
$\neg$ NAE: Denotes that we train a DeformTime model without the NAE module. For an input $\mathbf{Z}\!\in\!\mathbb{R}^{L\times(C+1)}$, we simply use a fully connected layer to embed the $C\!+\!1$ variables into a hidden dimension of size $d$ and obtain $\mathbf{E}\!\in\!\mathbb{R}^{L\times d}$.
$\neg\leavevmode\nobreak\ \mathbf{P}_{\texttt{n}}$: Denotes that we do not use the fixed position embedding $\mathbf{P}_{\texttt{n}}$ from the embedding procedure, i.e. instead of Equation 2, we use $\mathbf{Z}_{\texttt{e}}=\texttt{LN}\!\left(\mathbf{E}\right)$.
D.5 Computational complexity and efficiency of DeformTime – Additional information
The total number of operations (based on multiplications) for DeformTime is given by
$\underbrace{\left(C\!+\!1\right)dL\!+\!d}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small NAE}}}+2\!\underbrace{\left\{\underbrace{n\!\left[(k^{2}\!+\!1)\ell d\!+\!3\ell d^{2}\!+\!3\ell^{2}d\!+\!n\ell dL\right]}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small V-DAB}}}+\underbrace{r\!\left[(k\!+\!1)\kappa d\!+\!3\kappa d^{2}\!+\!3\kappa^{2}d\right]}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small T-DAB}}}+2d\!+\!6d^{2}L\right\}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small Encoder}}}+\underbrace{2d^{2}L}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small Decoder}}}\,.$ (13)
The main components that affect the order of operations are the following:
V-DAB: $\mathcal{V}=n\!\left[\left(k^{2}\!+\!1\right)\!\ell d+3\ell d^{2}+3\ell^{2}d+n\ell dL\right]\,\,\text{operations}$ (14)
$L\approx n\ell\implies\mathcal{O}\!\left(d^{2}L+dL^{2}\right)$
T-DAB: $\mathcal{T}=r\!\left[\left(k\!+\!1\right)\kappa d+3\kappa d^{2}+3\kappa^{2}d\right]\,\,\text{operations}$ (15)
$L=\kappa r\implies\mathcal{O}\!\left(d^{2}L\right)$
Encoder: $\mathcal{E}=2\!\left(\mathcal{V}+\mathcal{T}+2d+6d^{2}L\right)\,\,\text{operations}$ (16)
$\implies\mathcal{O}\!\left(d^{2}L+dL^{2}\right)$
Decoder: $\mathcal{G}=2d^{2}L\,\,\text{operations}$ (17)
$\implies\mathcal{O}\!\left(d^{2}L\right)\,.$
Hence, we conclude that the order of operations for DeformTime is $\mathcal{O}\!\left(d^{2}L+dL^{2}\right)$, where $L$ denotes the length (time steps) of the look-back window, and $d$ is the size of the hidden layers used throughout our method.
Figure D2: GPU memory (VRAM) consumption based on: (A) the look-back window ($L$), and (B) the number of input variables ($C\!+\!1$).
Figure D2 depicts the GPU VRAM memory consumption of DeformTime during training compared to other transformer-based models w.r.t. the length of the look-back window ($L$; part A) and the number of input variables ($C+1$; part B). In these experiments, we set both the batch size and the hidden dimension $d$ to $64$. While assessing the impact of $L$, we set $C\!+\!1$ to $32$, and while assessing the impact of the amount of input variables, we set $L$ to $56$. We then present the average memory consumption across $50$ training epochs.
Overall, DeformTime has a relatively small GPU memory footprint. As the length of the look-back window increases (from $28$ to $140$ time steps), iTransformer (lowest) and DeformTime (second lowest) are showing a stable memory consumption that does not exceed 111\!1 GB. We note that the low memory consumption of iTransformer is expected as it encodes the input over the sequence dimension, i.e. from $\mathbb{R}^{L\times(C+1)}$ to $\mathbb{R}^{d\times(C+1)}$, which guarantees consistent GPU memory usage regardless of the value of $L$. In DeformTime, the input partitioning and adaptive transformations within the V-DAB and T-DAB modules reduce computational complexity, leading to reduced memory consumption, especially when $L$ is large. At the same time, the memory footprints of PatchTST and Crossformer linearly increase with $L$.
As we now increase the number of variables from ($32$ to $512$), other methods display an exponential increase in memory consumption (delayed and considerably lower for iTransformer compared to PatchTST and Crossformer), whereas for DeformTime memory consumption is almost unaffected and stable (e.g. $<\!1\!$ GB for DeformTime, but $>\!24\!$ GB for Crossformer). Hence, we argue that our method can handle a greater amount of exogenous predictors more efficiently.
D.6 A brief note about the performance and evaluation of DLinear
Although DLinear achieved competitive results in prior work [53], the performance drop in our paper originates from the fact that we evaluate models based on their estimate for $y_{t+H}$, i.e. the target variable’s value at time step $t\!+\!H$ (see a detailed description of the forecasting task in section 2). In fact, for most tasks this is the target forecast (represented by the corresponding forecasting horizon $H$) and that is what forecasting accuracy should be measured on. Contrary to common sense, DLinear was evaluated on the entire time series (entire sequence of predictions), i.e. from $y_{t+1}$ to $y_{t+H}$. Oddly, this evaluation was conducted in a uniform way, i.e. the error for each time step incurred the same penalty. That may be relevant for some tasks, but it is not relevant to tasks where forecasting $H$ time steps ahead really necessitates obtaining an accurate forecast $H$ time steps ahead (and that should be the case for most, if not all, forecasting tasks). DLinear was obviously more accurate in early time steps (lower degree of difficulty), but very inaccurate in later ones (greater degree of difficulty). Hence, in our experiments DLinear displayed the worst forecasting performance.
D.7 Should MTS forecasting models mix variables in the embedding space?
The paper that presented the PatchTST method [32] provided some empirical evidence for the potential benefits of not modelling inter-variable dependencies. Contrary to that, our results demonstrate that this may not always hold for better models as well as different time series tasks and data sets. Based on our ILI rate forecasting experiments, we deduce that models that mix variables in the embedding space (Crossformer, LightTS, TimeMixer, iTransformer, and DeformTime) outperform the rest. Hence, and perhaps as expected, modelling inter-variable relationships should still be considered a viable approach in time series forecasting.
Appendix E Limitations
In this section, we summarise the potential limitations of this work. As with any machine learning model, the predictive capacity of a forecasting regression function is predominantly determined based on empirical evaluations. To this end, we have conducted extensive experiments. Nevertheless, there are undoubtedly more forecasting tasks that one could use to assess the performance of DeformTime. We have argued about the existence of potential issues within the currently deployed data sets in the evaluation of deep learning forecasters and we have proposed the use of new ones in this paper (alongside some of the previously established benchmarks). Although we hope that we will be able to share this data set with the broader community, at the time of submission we could only point to the application form for gaining access to the Google Health Trends API that we obtained the data from.
From a qualitative analysis standpoint, it would be interesting to provide further insights about DeformTime’s performance under a controlled amount of exogenous predictors. We do make the claim that DeformTime is designed to handle exogenous predictors better and experiment with new data sets that contain significantly more exogenous predictors compared to the established benchmark data sets (and show increased performance gains). However, we do not assess this under a more rigorous experimental design. We could add that our ablation study offers some level of interpretation for the contributions of different components. Nonetheless, further experiments would be required to more concretely understand DeformTime’s modelling strengths and weaknesses.
From a methodological perspective, there might be ways to further improve DeformTime. These, of course, could be considered as future avenues of research (as opposed to limitations). First, the deformations we are performing are localised. Perhaps a more expansive deformation approach would be capable of capturing more intricate relationships that a correlation-based neighbourhood omits. In addition, and this applies to all deep learning MTS models, DeformTime does not offer any confidence bounds for its forecasting estimates. We could arguably achieve that via conformal prediction or other approaches. However, how would that affect the accuracy of our mean estimates and the overall model complexity? This is certainly an avenue for follow up work, especially if a forecasting model is to be deployed in practical applications.
Appendix F Societal impact
As with any scientific development, our work may have positive but also negative societal impact. From a positive perspective, forecasting models can be used to better predict outcomes in various domains and improve quality of life. Contrary to that, forecasting models could also be exploited by malicious actors in various forms of malpractice. We do think this should be common knowledge. We do also note that there is a thin line between safeguarding AI malpractice and significantly restricting scientific progress (and hence, quality of life improvements). This is beyond the scope of our work and is obviously a broader topic of discussion.
We would like to note that outcomes of this work (e.g. an improved variant of DeformTime) may be used in infectious disease monitoring systems by interested stakeholders (e.g. public health organisations such as CDC, ECDC, UKHSA etc.). These, of course, do not constitute malicious actors. However, a potential risk may arise when (if) a forecasting model provides inaccurate insights. Nevertheless, this is a shared risk among various other disease models (including established mechanistic models). To mitigate this risk, public health agencies use multiple endpoints to determine their course of action (e.g. sentinel and syndromic surveillance, laboratory tests, rapid tests, and vetted epidemiological models). Hence, any decision is based on a collection of different predictors. Nonetheless, prior to the adoption of any forecasting model within public policy, a more thorough and focused (to the said application domain) evaluation of the forecasting method is required.
Appendix G Supplementary ILI rate forecasting figures
Here we present a series of figures for the ILI rate forecasting task. Each figure shows the estimates from all models for a certain forecasting horizon and location.
Figure G3: 7 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure G4: 14 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure G5: 21 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure G6: 28 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure G7: 7 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure G8: 14 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure G9: 21 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure G10: 28 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure G11: 7 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).
Figure G12: 14 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).
Figure G13: 21 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).
Figure G14: 28 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).